OpenStudy (anonymous):

10 math questions !! Help please!!! I WOULD APPRECIATE IT SO MUCH! (you dont have to do all of them just as much as you can) Medal will be awarded for the person who helps the most!

OpenStudy (anonymous):

1. –2v – 7 = –23 (1 point) 15 8 –8 –15 2. x over three – 10 = –12 (1 point) –6 –66 6 66 3. x over four + 10 = 14 (1 point) 96 46 16 –16 4. 3(4 – 2x) = –2x (1 point) –1 1 2 3 5. 3.75x + 3.7 = 1.7 + 1.75x (1 point) 10 –1 1 .1 For questions 6–7, simplify the expression. 6. –9 – 6(w + 5) (1 point) –15w + 5 –15w + 30 –6w – 39 –6w + 21 7. 2x + 3(x – 2) – 3(x – 6) (1 point) 2x + 12 2x – 8 8x – 24 2x – 24 For questions 8–10, solve the inequality. 8. c + 6 < –20 (1 point) c < –26 c < –14 c < 26 c < 14 9. d over thirteen ≥ –3 (1 point) d ≥ –16 d ≥ 39 d ≥ 10 d ≥ –39 10. –8x > –64 (1 point) x < 8 x > 8 x < –72 x > 72

OpenStudy (anonymous):

lol